# USE-LFA: A Data-Driven Framework for Urban Air Mobility Site Evaluation using Latent Factor Analysis

### v1

@@ Abstract
This study introduces the Urban Site Evaluation using Latent Factor Analysis (USE-LFA) framework, a data-driven methodology for selecting optimal urban port locations, demonstrated through a case study in Seoul. USE-LFA identifies six critical urban factors categorized into suitability (operational feasibility) and attractiveness (operational merits), enabling a holistic evaluation of potential sites. The framework employs latent factor analysis to uncover hidden patterns in urban data, integrating diverse attributes such as floating population, economic activity, and infrastructure connectivity. A composite metric, the v-score, balances suitability and attractiveness, offering flexibility to prioritize specific strategic goals. Spatial analysis reveals distinct regional typologies—balanced, suitability-biased, and attractiveness-biased regions—highlighting the adaptability of USE-LFA across varying urban contexts. Beyond its application to vertiports, this versatile framework provides valuable insights for integrating emerging mobility technologies and transit hubs into complex urban environments, supporting sustainable urban planning and informed policymaking.
@@ 1 Introduction
Existing studies on vertiport site selection have employed various methodologies such as clustering algorithms, simulation-based methods, custom optimization models, expert surveys, and GIS-based approaches. These approaches have provided valuable insights and laid a strong foundation for understanding this complex problem. Nevertheless, given the intricate dynamics of urban environments, there remains an opportunity to expand upon these efforts by incorporating more comprehensive perspectives on urban structures and dynamics.
To contribute to this growing body of work, we propose a novel framework leveraging Latent Factor Analysis (LFA) on regional data to uncover latent factors that characterize the urban environment for vertiport integration. Our approach introduces the concept of location-signatures, the structural latent factors that constitute urban environments, grouped into two pillars: suitability and attractiveness.
@@
LFA emerges as a critical solution, uncovering and quantifying latent location-signatures that reflect underlying urban structures and dynamics. This approach enables a comprehensive evaluation that integrates social considerations like community dynamics with economic factors such as economic activeness. By consolidating these various factors into a composite v-score metric, LFA offers a holistic perspective on urban environments, ensuring sustainable vertiport integration that balances operational feasibility with operational merits.
This paper is structured as follows: Section 2 provides details on existing work with diverse methodologies. Section 3 introduces the framework and methodology. Section 4 presents data and results for Seoul case study, and Section 5 discussion furhter result analysis. Finally, Section 6 concludes the paper with suggestions for future studies.
@@ 3 Methodology
We introduce USE-LFA, a novel framework for UAM site evaluation framework using latent factor analysis (LFA) that encapsulates multiple urban considerations into suitability and attractiveness. USE-LFA is employed on regional data to identify latent factors that characterize urban environments relevant to vertiport integration (see Figure 1). This study employs principal axis factoring (PAF) with Varimax rotation and regression-based scoring to identify latent factors underlying urban attributes across regions.
Figure 1: Flow of VESLA, from the city-specific urban attributes to the suitability and attractiveness scores for the regions.
3.1 Input Data Preparation for VESLA
The table below provides an example of data attributes that can be used as input for VESLA. The vertiport-related themes are predetermined and relevant data is gathered accordingly. The input attributes are flexible and can be tailored to suit the unique characteristics of each study area. The description column outlines the normalization process, an essential step for LFA to address variations in census tract sizes and population densities across the regions.
@@ 3.3 Latent Factor Analysis (LFA)
The core constitution of VESLA is a robust mathematical model that decomposes complex regional data into interpretable factors. Let $\mathbf{A}\in\mathbb{R}^{N\times R}$ represent the standardized urban attribute matrix where $N$ is the number of standardized urban attributes (indexed by $i=1,2,\dots,N$) and $R$ is the number of regions in the study area (indexed by $j=1,2,\dots,R$). Each element $a_{ij}$ encodes the standardized measurement of urban attribute $i$ in region $j$.
The foundational relationship between the urban attributes and their underlying latent factors is formalized as:
@@
The framework distills complex urban attributes into interpretable dimensions by employing factor analysis to identify latent urban factors, and then strategically categorizing these factors into distinct Suitability and Attractiveness. This binary categorization offers a structured analytics of a dual-perspective evaluation of urban regions while acknowledging the multidimensional nature of urban environment. The framework’s mathematical formalization ensures reproducibility across different urban contexts, while its flexibility allows researchers and planners to adapt factor classifications based on specific policy objectives or theoretical considerations.
@@ 4.4 Suitability and Attractiveness Scores
The framework categorizes urban factors into two dimensions critical for UAM site selection: suitability (operational feasibility) and attractiveness (operational merits). Figure 3 illustrates their spatial distribution across Seoul, revealing distinct geographic patterns. Central business districts exhibit concentrated attractiveness scores, reflecting their inherent advantages in demand generation and economic activity. In contrast, suitability scores display broader dispersion, with elevated values observed in less congested peripheral neighborhoods. This dichotomy establishes a foundational tension between immediate operational viability and long-term scalability in vertiport planning.
@@ 6 Conclusion
Unlike existing tools such as NASA’s VAMOS! (Sheth, 2023), which often conflate operational feasibility and merits into a single evaluation, USE-LFA distinctly separates these dimensions while integrating broader urban environmental factors such as community preparedness and societal equity. This separation enhances the framework’s versatility, allowing it to adapt to diverse urban contexts and enabling stakeholders to evaluate suitability and attractiveness independently while tailoring decisions to specific regional needs. Furthermore, its flexible design supports applications beyond vertiports, such as transit hubs and other emerging mobility infrastructures. By systematically addressing the interplay between operational feasibility and metrits, USE-LFA provides informative insights for the systematic integration of ports into complex urban environments. Policymakers can leverage this framework to align port placement with strategic goals such as improving accessibility in underserved areas or enhancing emergency response capabilities. Urban planners can use factor-specific insights to design infrastructure that complements existing systems while addressing community concerns. Moreover, the framework’s flexibility allows for continuous adaptation to evolving urban environments or new data inputs.

### v2

@@ Abstract
Urban air mobility (UAM) introduces new challenges for infrastructure planning, requiring data-driven approaches for sustainable site selection. This study proposes USE-LFA (Urban Site Evaluation using Latent Factor Analysis), a framework designed to support equitable and environmentally conscious siting of urban ports. Applying latent factor analysis to 25 urban attributes in Seoul, the framework identifies six latent factors, grouped into two dimensions: Suitability and Attractiveness. These dimensions are combined through a tunable prioritization metric, enabling alignment with local strategic goals. The analysis uncovers spatial typologies and clustered siting patterns, highlighting regional disparities in site potential. Sensitivity analysis demonstrates that small adjustments in the Suitability–Attractiveness weighting substantially affect viable site candidates, emphasizing the need for calibrated decision-making. USE-LFA facilitates interpretable and transferable analysis across different urban contexts and datasets, offering a scalable approach to integrating UAM and other emerging mobility systems into urban environments, while advancing sustainable and inclusive transport infrastructure development.
@@ 1 Introduction
Existing studies on vertiport site selection have employed various methodologies such as clustering algorithms, simulation-based methods, custom optimization models, expert surveys, and GIS-based approaches. These approaches have provided valuable insights and laid a strong foundation for understanding this complex problem. Nevertheless, given the intricate dynamics of urban environments, there remains an opportunity to expand upon these efforts by incorporating more comprehensive perspectives on urban environments.
To contribute to this growing body of work, we propose a novel framework leveraging Latent Factor Analysis (LFA) on regional data to uncover latent factors that characterize the urban environment for vertiport integration. These latent factors are grouped into two pillars: suitability and attractiveness.
@@
LFA emerges as a critical solution, uncovering and quantifying latent factors that reflect underlying urban structures and dynamics. This approach enables a comprehensive evaluation that integrates social considerations like community preparedness with economic factors such as economic dynamism. By consolidating these various factors into a composite v-score metric, LFA offers a holistic perspective on urban environments, ensuring sustainable vertiport integration that balances operational feasibility with operational merits.
This paper is structured as follows: Section 2 provides details on existing work with diverse methodologies. Section 3 introduces the framework and methodology. Section 4 presents data and results for Seoul case study, and Section 5 discussion further result analysis. Finally, Section 6 concludes the paper with suggestions for future studies.
@@ 3 Methodology
We propose the Urban Site Evaluation using Latent Factor Analysis (USE-LFA) framework, a data-driven approach that captures the multifaceted nature of urban environments by disentangling them into two distinct dimensions: suitability and attractiveness. USE-LFA applies latent factor analysis (LFA) to regional urban data to uncover latent structures that characterize urban environments relevant to vertiport integration (see Figure 1). This study employs principal axis factoring (PAF) with Varimax rotation and regression-based scoring to identify latent factors underlying urban attributes across regions.
Figure 1: Flow of USE-LFA, from the city-specific urban attributes to the suitability and attractiveness scores for each region.
3.1 Input Data Preparation for USE-LFA
The table below provides an example of data attributes that can be used as input for USE-LFA. The vertiport-related themes are predetermined and relevant data is gathered accordingly. The input attributes are flexible and can be tailored to suit the unique characteristics of each study area. The description column outlines the normalization process, an essential step for LFA to address variations in census tract sizes and population densities across the regions.
@@ 3.3 Latent Factor Analysis (LFA)
The core constitution of USE-LFA is a robust mathematical model that decomposes complex regional data into interpretable factors. Let $\mathbf{A}\in\mathbb{R}^{N\times R}$ represent the standardized urban attribute matrix where $N$ is the number of standardized urban attributes (indexed by $i=1,2,\dots,N$) and $R$ is the number of regions in the study area (indexed by $j=1,2,\dots,R$). Each element $a_{ij}$ encodes the standardized measurement of urban attribute $i$ in region $j$. The foundational relationship between the urban attributes and their underlying latent factors is formalized as:
@@
The framework distills complex urban attributes into interpretable dimensions by employing latent factor analysis to identify critical urban factors, and then strategically categorizing these factors into Suitability and Attractiveness. This binary categorization offers a structured analytics of a dual-perspective evaluation of urban regions while acknowledging the multidimensional nature of urban environment. The framework’s mathematical formalization ensures reproducibility across different urban contexts, while its flexibility allows researchers and planners to adapt factor classifications based on specific policy objectives or theoretical considerations.
@@ 4.4 Suitability and Attractiveness Scores
The framework categorizes urban factors into two dimensions critical for UAM site selection: suitability (operational feasibility) and attractiveness (operational merits). Figure 3 illustrates their spatial distribution across Seoul, revealing distinct geographic patterns. Central business districts exhibit concentrated attractiveness scores, reflecting their inherent advantages in demand generation and economic activity. In contrast, suitability scores display broader dispersion, with elevated values observed in less congested peripheral regions. This dichotomy establishes a foundational tension between immediate operational viability and long-term scalability in vertiport planning.
@@ 6 Conclusion
Unlike existing tools such as NASA’s VAMOS! (Sheth, 2023), which often conflate operational feasibility and merits into a single evaluation, USE-LFA distinctly separates these dimensions while integrating broader urban environmental factors such as community preparedness and societal equity. This separation enhances the framework’s versatility, allowing it to adapt to diverse urban contexts and enabling stakeholders to evaluate suitability and attractiveness independently while tailoring decisions to specific regional needs. Furthermore, its flexible design supports applications beyond vertiports, such as transit hubs and other emerging mobility infrastructures. By systematically addressing the interplay between operational feasibility and merits, USE-LFA provides informative insights for the systematic integration of ports into complex urban environments. Policymakers can leverage this framework to align port placement with strategic goals such as improving accessibility in under-served areas or enhancing emergency response capabilities. Urban planners can use factor-specific insights to design infrastructure that complements existing systems while addressing community concerns. Moreover, the framework’s flexibility allows for continuous adaptation to evolving urban environments or new data inputs.